### v1

# Seer: Facilitating Structured Reasoning and Explanation via Reinforcement Learning

Abstract
Elucidating the reasoning process with structured explanations from question to answer is fundamentally crucial, as it significantly enhances the interpretability and trustworthiness of question-answering (QA) systems. However, structured explanations demand models to perform intricate structured reasoning, which poses great challenges. Most existing methods focus on single-step reasoning through supervised learning, ignoring logical dependencies between steps. Meanwhile, existing reinforcement learning (RL)-based methods overlook the structured relationships, impeding RL’s potential in structured reasoning. In this paper, we propose Seer, a novel method that maximizes a structure-based return to facilitate structured reasoning and explanation. Our proposed structure-based return precisely describes the hierarchical and branching structure inherent in structured reasoning, effectively capturing the intricate relationships between states. We also introduce a fine-grained reward function to meticulously delineate diverse reasoning steps. Extensive experiments show that Seer significantly outperforms state-of-the-art methods, achieving an absolute improvement of 6.9% over RL-based methods on EntailmentBank, a 4.4% average improvement on STREET benchmark, and exhibiting outstanding efficiency and cross-dataset generalization performance.
1 Introduction
Navigating machines to reason and explain the reasoning process from question to answer is a long-term pursuit in the AI community (McCarthy, 1959; Yu et al., 2023). While existing QA systems adeptly furnish brief supporting evidence (Rajani et al., 2019; DeYoung et al., 2020), they often fail to clarify the reasoning process from prior knowledge to the answer. This means that it is difficult to understand how the answer is reached. However, by elucidating the reasoning process of predicted answers, we can observably improve interpretability, trustworthiness, and debuggability in QA systems (Dalvi et al., 2021; Ribeiro et al., 2023). As illustrated in Figure 1, when answering the question "Which natural material is best for making a table?", the reasoning process with structured explanations, such as entailment trees (Dalvi et al., 2021) or reasoning graphs (Ribeiro et al., 2023), can explain why "sturdy wood" is the predicted answer, which in turn increases trustworthiness.
Figure 1: An example of structured explanation. Given a hypothesis $h$ (a declarative sentence derived from a question-answer pair) and a set of facts (or corpus), the goal is to generate a structured explanation, which delineates the reasoning process from facts to the hypothesis.
Method Reasoning Type Runtime Return
RLET multi-step reasoning 9.34s chained
FAME single-step reasoning 30.77s /
Ours structured reasoning 3.91s structured
Table 1: Comparative analysis of different methods: RL-based method, RLET (Liu et al., 2022), supervised method, FAME (Hong et al., 2023), and our approach.
However, deriving those complex structured explanations poses a great challenge. Previous methods (Dalvi et al., 2021; Tafjord et al., 2021) consider them as linearized sequences and generate the entire reasoning process at once. However, these methods lack controllability and may hallucinate unreliable reasoning steps. To address these concerns, recent studies (Hong et al., 2022; Neves Ribeiro et al., 2022; Yang et al., 2022) decompose structured explanations and focus on single-step reasoning via supervised learning. Nevertheless, this approach may not always yield optimal results as it fails to consider the interdependencies between different steps. FAME (Hong et al., 2023) attempts to compensate for these shortcomings by leveraging Monte-Carlo planning (Kocsis and Szepesvári, 2006). Yet, as illustrated in Table 1, it significantly increases the running time and inadvertently explores numerous ineffective steps. Furthermore, FAME still concentrates on isolated single-step reasoning, which hardly facilitates multi-step reasoning. Conversely, RLET (Liu et al., 2022) endeavors to enhance the multi-step reasoning through RL. However, RLET still defines the return (a.k.a cumulative rewards) in a chained trajectory, thereby overlooking the inherent tree or graph logical structure in structured reasoning (Dalvi et al., 2021; Ribeiro et al., 2023). Such simplifications may impede the potential of RL for structured reasoning. For instance, RLET may allocate different returns for equivalent trajectories (Yu et al., 2023).
To address the above issues, we propose Seer, a novel method that facilitates Structured rEasoning and Explanation via Reinforcement learning. In structured reasoning, we observe that the logical dependencies between different steps no longer follow a chained trajectory but instead adhere to the inherent tree or graph structure. Therefore, we propose the structure-based return to precisely describe a tree or graph logical structure, effectively capturing the complex interdependencies between different steps. Furthermore, we refine the reward function to meticulously delineate diverse reasoning steps, particularly redundant ones, which do not play a role in the final structured explanations. Through experiments in Sec. 5.4, we find that redundant steps represent the exploration in the environment, and appropriate penalization contributes to improved reasoning performance.
Our contributions are summarized as follows:
$\bullet$ We propose Seer, a novel RL-based method that facilitates structured reasoning and explanation. To our knowledge, Seer is the first general framework that accommodates scenarios of chained, tree-based, and graph-based structured reasoning.
$\bullet$ We also propose the structure-based return to address the intricate interdependencies among different reasoning steps, effectively stimulating the potential of RL in structured reasoning.
$\bullet$ Extensive experiments demonstrate that Seer significantly outperforms state-of-the-art methods. It not only facilitates the effectiveness and efficiency in structured reasoning but also exhibits outstanding cross-dataset generalization performance.
2 Related Work
2.1 Explanation for Question Answering
Extensive research has delved into various forms of explainability in QA systems (Thayaparan et al., 2020; Wiegreffe and Marasovic, 2021; Lamm et al., 2021). Different from the free-form texts susceptible to hallucinations (Rajani et al., 2019; Wei et al., 2022) or the rationales that only provide supporting evidence (DeYoung et al., 2020; Valentino et al., 2021), the structured explanations, such as the entailment trees (Dalvi et al., 2021) and reasoning graphs (Ribeiro et al., 2023), offer a novel way to generate explanations with structured reasoning in a tree or graph structure, delineating what and how knowledge can be integrated to reach the answer. Despite the remarkable interpretability, the intricate structured reasoning also poses significant challenges (Yu et al., 2023; Xu et al., 2023).
2.2 Natural Language Reasoning
Natural language reasoning, a process that integrates multiple knowledge to derive new conclusions, has attracted significant attention (Saha et al., 2020; Tafjord et al., 2021; Sanyal et al., 2022). Among these, the entailment trees and reasoning graphs, which involve structured reasoning and reasoning path generation tasks, present considerable challenges (Yu et al., 2023). Dalvi et al. (2021) attempts to transform the structured reasoning into a linearized sequence to fit encoder-decoder models, which may generate hallucinations and invalid reasoning. To alleviate this issue, recent studies (Neves Ribeiro et al., 2022; Hong et al., 2022; Neves Ribeiro et al., 2022; Hong et al., 2023) perform premises selection and reasoning in a step-by-step manner. Nevertheless, these methods decompose structured reasoning and solely leverage isolated single-step supervision to train models. This approach neglects the interdependencies between steps, which may not always yield optimal results. Therefore, in light of the advancements of RL in various reasoning tasks (Poesia et al., 2021; Le et al., 2022), RLET (Liu et al., 2022) attempts to incorporate RL into the entailment trees. However, it has to enumerate all potential actions, which is evidently unacceptable for practical scenarios. Furthermore, RLET still defines returns in chained trajectories to facilitate multi-step reasoning, which is obviously not suitable for tree/graph-based structured reasoning. In contrast, our Seer showcases superior adaptability to chained, tree-based, and graph-based structured reasoning via the structure-based return, which significantly enhances both the reasoning performance and efficiency.
3 Method
3.1 Task Formulation
As illustrated in Figure 1, the input of the task comprises a set of facts $X=\{x_{1},x_{2},\ldots,x_{n}\}$ and a hypothesis $h$. The output of the task is an entailment tree $T$ (or a reasoning graph) [^1] For the sake of brevity, we formalize the task using the entailment tree (Dalvi et al., 2021) as an example, and the reasoning graph (Ribeiro et al., 2023) is similar.. The entailment tree $T$ consists of tree-structured reasoning, whose leaf nodes are selected from the relevant facts ($x_{*}$) and intermediate nodes represent the derived intermediate conclusions ($i_{*}$). We represent the annotated groundtruth entailment tree as $T_{gold}$, with its leaf nodes signifying $X_{gold}$.
Figure 2: Overall framework of our Seer.
3.2 Overview
Figure 2 illustrates the overall framework of our proposed Seer, which mainly includes trajectory collection and parameter updates. For trajectory collection, we delineate the state update process into two steps: action generation and conclusion reasoning. The trajectory is produced iteratively until the stopping condition is satisfied (Appendix C.1). For parameter updates, the trajectories are converted to $T_{\text{pred}}$ and aligned with intermediate nodes in $T_{\text{gold}}$ based on maximum Jaccard similarity. Please refer to Appendix C.2 for more alignment details. Then, the policy and critic are updated based on the structure-based return. Algorithm 1 outlines our proposed method for further reference.
3.3 Fine-grained Component of Seer
State
We define the state $s_{t}=\{h,P_{t},C_{t}\}$ as a combination of the hypothesis $h$, existing reasoning steps $P_{t}$ and candidate sentences $C_{t}$. $P_{t}$ contains the reasoning steps so far. $C_{t}$ is the set of sentences that can be selected as premises. Each sentence in $C_{t}$ is either unused facts or intermediate conclusions $I_{t}$ generated by previous steps, i.e., $C_{t}=\{X\cup I_{t}\backslash U_{t}\}$, where $U_{t}$ is the set of used sentences. For the initial state, $s_{1}=\{h,P_{1}=\varnothing,C_{1}=X\}$. Evidently, the existing reasoning steps $P_{t}$ and candidate sentences $C_{t}$ entirely depend on the previous state $s_{t-1}$ and its actions $a_{t-1}$. Thus, the state updates process adheres to the Markov decision process (Bellman, 1957).
Action
[…]
Policy
While the action "Reason: <premises>" facilitates reasoning with combinations of premises, it also significantly increases the complexity of the action space. This is because <premises> can be any combination of elements from the candidate set $C$. Previous studies (Liu et al., 2022; Hong et al., 2022) opted to limit combinations to pairwise premises, resulting in $\binom{n}{2}$ potential actions, where $n$ is the size of the set $C$. However, this approach has apparent limitations. Firstly, as candidate sentences $C$ increases, the number of enumerated actions will experience exponential growth. This makes it computationally expensive and impractical for larger problems. Secondly, by restricting combinations to pairs only, it potentially overlooks the interdependencies among multiple premises, which may limit the effectiveness and richness of the derived conclusions.
To address this issue, we adopt a generative model (Raffel et al., 2020) as the policy $\pi$ to effectively manage the action space $\mathcal{A}$. Despite the vastness of the theoretical action space deriving from "Reason: <premises>", we observe that the number of effective actions in practice is considerably smaller. Therefore, instead of enumerating all possible pairs which can be computationally expensive and may overlook complex interdependencies, our policy $\pi$ samples from the current action space to generate the top-$k$ effective actions.
$a_{t}^{1},a_{t}^{2},...,a_{t}^{k}\sim\pi(a|s_{t}),\quad a\in\mathcal{A}(s_{t})$ (1)
where the input is a linearized state $s_{t}$, which is the concatenation of $h$, $P_{t}$, and $C_{t}$. This not only reduces the computational complexity but also effectively manages the premises with complex combinations, rather than being confined to paired premises. Furthermore, it is important to note that the probabilities of these actions may not sum up to 1, as they are a subset of the entire action space. Therefore, we proceed with re-normalization to form an appropriate probability distribution over the top-$k$ actions, and sample from it to select the current action $a_{t}$.
$\pi^{\prime}(a_{t}^{i}|s_{t})=\frac{\pi(a_{t}^{i}|s_{t})}{\sum_{j=1}^{k}\pi(a_{t}^{j}|s_{t})},\quad i=1,...,k$ (2)
$a_{t}\sim\pi^{\prime}(a|s_{t}),\quad a\in\{a_{t}^{1},a_{t}^{2},...,a_{t}^{k}\}$ (3)
Consequently, this policy allows us to explore the action space more efficiently and effectively.
Entailment Module
If the action $a_{t}$ is "Reason: <premises>", we invoke the entailment module to derive the intermediate conclusion to obtain the next state. The entailment module is also a generative model with its input being <premises>. Following (Hong et al., 2022; Liu et al., 2022), we fine-tune the entailment model in a supervised manner and freeze the parameters during the reinforcement learning process, as depicted in Figure 2.
Reward
To evaluate the correctness of the entailment tree, Dalvi et al. (2021) proposed an alignment algorithm based on Jaccard similarity to align each intermediate node of the predicted tree $T_{\text{pred}}$ with $T_{\text{gold}}$. However, different from the fully supervised learning methods, we observe that during the RL process, the policy may explore some redundant steps, i.e., these steps do not contribute to the reasoning for the final hypothesis. Existing RL-based work (Liu et al., 2022) simply treats redundant steps with the same penalty as erroneous steps. This simplification may adversely affect the learning process. Firstly, it may discourage necessary exploration in the action space. Secondly, it does not provide meticulous feedback to guide the policy towards optimal policy, since it fails to distinguish between innocuous actions (redundant steps) and incorrect actions (erroneous steps).
To address this issue, we propose a fine-grained reward function that takes into account correct steps, erroneous steps, and redundant steps. For a trajectory $\tau$, we assume that the last intermediate conclusion is our predicted hypothesis, due to the policy deems it should End here. Then, based on the entailment relationships between premises and intermediate conclusions, we backtrack to construct the predicted entailment tree $T_{\text{pred}}$. Note that there might be some steps not participating in the entailment tree, which are regarded as redundant steps. We consider steps that perfectly match via the alignment algorithm (Dalvi et al., 2021) as correct steps and regard others as erroneous steps.
$r_{t}=\begin{cases}1,&\text{if }\text{perfectly match},\\ -0.5,&\text{if }i_{t}\notin T_{\text{pred}},\\ -1,&\text{otherwise}.\end{cases}$ (4)
Critic
To enhance training stability, we introduce the critic that estimates the state-value function $V(s_{t})$. The input of $V(s_{t})$ is a linearized state, and the output is a scalar representing the return starting from state $s_{t}$. Then, we employ one-step temporal difference strategy (Sutton, 1988) to incrementally adjust these estimated values:
$G_{t}=r_{t}+\gamma V(s_{t+1})$ (5)
where $\gamma$ is the discount factor. Equation 5 brightly describes the chained multi-step reasoning. However, in structured reasoning, reasoning steps typically adhere to inherent tree or graph structures (Dalvi et al., 2021; Ribeiro et al., 2023). The chained structure is merely a special case among them. Thus, equation 5 may not effectively capture the intricate logical dependencies between steps in structured reasoning.
Therefore, we propose the structure-based return to accommodate the logical dependencies inherent in such structured reasoning.
$\hat{G}_{t}=r_{t}+\gamma\frac{1}{|\mathcal{P}(s_{t})|}\sum_{s_{j}\in\mathcal{P}(s_{t})}V(s_{j})$ (6)
where $\mathcal{P}(s_{t})$ represents the parent node of state $s_{t}$ in the entailment tree $T_{\text{pred}}$ or reasoning graph. When $s_{t}\notin T_{\text{pred}}$, $\mathcal{P}(s_{t})=s_{t+1}$. It is worth noting that in the entailment tree, each state typically has only one parent node, i.e., $\hat{G}_{t}=r_{t}+\gamma V(\mathcal{P}(s_{t}))$. Therefore, the entailment tree is also a special case of reasoning graphs (Ribeiro et al., 2023). However, our structure-based return (equation 6) can comprehensively adapt to structured reasoning involving chained, tree-based and graph-based structured scenarios. For example, as shown in Figure 6, for equivalent trajectories $s_{1}\rightarrow s_{2}\rightarrow s_{3}\rightarrow s_{4}$ and $s_{2}\rightarrow s_{1}\rightarrow s_{3}\rightarrow s_{4}$, previous method (Liu et al., 2022) would assign different returns for state $s_{1}$ and $s_{2}$, even though they finally represent the same tree (Figure 6). Conversely, our method, by precisely delineating the intricate interdependencies between reasoning steps, consistently allocates the same return to any equivalent trajectories, thereby enhancing both stability and effectiveness.
3.4 Optimization
Our objective is to enhance structured reasoning capabilities of the policy through RL. To alleviate issues of training instability and sample inefficiency in RL (Zhou et al., 2023; Roit et al., 2023), we employ the proximal policy optimization (Schulman et al., 2017) as follows:
$\begin{gathered}\mathcal{L}_{\pi}=\mathbb{E}_{t}[\min\Big{(}\frac{\pi^{\prime}_{\theta}(a_{t}|s_{t})}{\pi^{\prime}_{\theta_{old}}(a_{t}|s_{t})}\hat{A}_{t},\text{clip}\big{(}\frac{\pi^{\prime}_{\theta}(a_{t}|s_{t})}{\pi^{\prime}_{\theta_{old}}(a_{t}|s_{t})},\\ 1-\epsilon,1+\epsilon\big{)}\Big{)}\hat{A}_{t}+\beta\mathcal{E}(\pi^{\prime}_{\theta})]\end{gathered}$ (7)
where $\pi^{\prime}$ represents the probabilities normalized by equation 2, $\theta$ and $\theta_{old}$ are parameters of the new and old policies, $\epsilon$ is a hyperparameter defining the clipping range, $\beta$ is the entropy exploration coefficient, and $\mathcal{E}$ is the entropy bonus, which encourages sufficient exploration.
$\mathcal{E}(\pi^{\prime}_{\theta})=\mathbb{E}_{a_{t}\sim\pi_{\theta}}[-\log\pi^{\prime}_{\theta}(a_{t}|s_{t})]$ (8)
Futhermore, $\hat{A}_{t}$ is the estimate of the advantage function for state $s_{t}$, defined as follows:
$\hat{A}_{t}=\hat{G}_{t}-V(s_{t})$ (9)
To accurately evaluate future returns and guide the policy towards better updates, we minimize the difference between the actual return and the predicted return by the critic.
$\mathcal{L}_{V}=\mathbb{E}_{t}\left[(V(s_{t})-\hat{G}_{t})^{2}\right]$ (10)
Supervised Warm-up
Incorporating the supervised warm-up strategy before RL offers a relatively stable initial policy, which facilitates faster adaptation to the environment, particularly for complex reasoning tasks (Ramamurthy et al., 2023; Wu et al., 2023). Therefore, we convert the structured reasoning into single-step supervised data to warm up the policy as follows:
$\mathcal{L}_{\text{warmup}}=-\sum_{i}\log p(y_{i}|s,y_{<i})$ (11)
where $s$ is the linearized state.
4 Experiments
4.1 Datasets
Tree-structured reasoning
We conduct experiments on EntailmentBank (Dalvi et al., 2021), the first dataset that supports structured explanation with entailment trees. Following (Hong et al., 2023), we also conduct experiments on EntailmentBankQA (Tafjord et al., 2022), whose objective is to reach the answer based on the entailment tree.
Graph-structured reasoning
We conduct experiments on the STREET benchmark (Ribeiro et al., 2023) to assess the performance of graph-structured reasoning. Please refer to Appendix B for more details about the dataset statistics.
4.2 Baselines
For EntailmentBank, we compare with single-pass methods, such as EntailmentWriter (Dalvi et al., 2021), and step-by-step methods including METGEN (Hong et al., 2022), IRGR (Neves Ribeiro et al., 2022), RLET (Liu et al., 2022), NLProofs (Yang et al., 2022), and FAME (Hong et al., 2023). For EntailmentBankQA, we compare with Selection-Inference (SI) (Creswell and Shanahan, 2022) and FAME (Hong et al., 2023). For the STREET benchmark, we compare with the method proposed in (Ribeiro et al., 2023). Furthermore, we conduct comparisons with GPT-4 (OpenAI, 2023) equipped with Chain-of-Thought (CoT) (Wei et al., 2022), Tree of Thought (ToT) (Yao et al., 2023a) and ReAct (Yao et al., 2023b).
4.3 Implementation Details
For a fair comparison [^2] Previous studies have consistently utilized T5-large as the base model. Despite the existence of more advanced generative models (Du et al., 2022; Touvron et al., 2023), using T5-large enables us to maintain a fair comparison., the policy is built with a T5-large model (Raffel et al., 2020), while the critic is the encoder of T5-large combined with a MLP ($\tanh$ as the activation function). For supervised warm-up, we set a learning rate of 1e-5, a batchsize of 16 and train the model for 20 epochs. For RL training, we set learning rate 2e-6 for both policy and critic, discounter factor $\gamma$ as 0.95, batchsize as 3, buffer size as 12, buffer training epochs $N_{K}$ as 2, $\epsilon$ as 0.2, and $\beta$ as 1e-4. More implementation details can be found in Appendix C
4.4 Evaluation Metrics
For EntailmentBank, we evaluate $T_{\text{pred}}$ with the following dimensions: Leaves, Steps, Intermediates and Overall AllCorrect. For STREET benchmark, we evaluate the reasoning graphs with two dimensions: Answer Accuracy and Reasoning Graph Accuracy. Note that Overall AllCorrect and Reasoning Graph Accuracy are extremely strict metrics, where any deviations will result in a score of 0. More metrics details can be found in Appendix D.
Task Method Leaves Steps Intermediates Overall
F1 AllCorrect F1 AllCorrect F1 AllCorrect AllCorrect
Task1 EntailmentWriter 98.7 84.1 50.0 38.5 67.6 35.9 34.4
METGEN 100.0 100.0 57.9 42.1 71.3 39.2 37.0
IRGR 97.6 89.4 50.2 36.8 62.1 31.8 32.4
RLET 100.0 100.0 54.6 40.7 66.9 36.3 34.8
NLProofS 97.8 90.1 55.6 42.3 72.4 40.6 38.9
Seer (Ours) 100.0 100.0 67.6 52.6 70.3 42.6 40.6
Task2 EntailmentWriter 83.2 35.0 39.5 24.7 62.2 28.2 23.2
METGEN 83.7 48.6 41.7 30.4 62.7 32.7 28.0
IRGR 69.9 23.8 30.5 22.3 47.7 26.5 21.8
RLET 81.0 39.0 38.5 28.4 56.3 28.6 25.7
NLProofS 90.3 58.8 47.2 34.4 70.2 37.8 33.3
Seer (Ours) 86.4 53.5 56.8 39.7 66.3 38.3 34.7
Task3 EntailmentWriter 35.7 2.9 6.1 2.4 33.4 7.7 2.4
METGEN 34.8 8.7 9.8 8.6 36.7 20.4 8.6
IRGR 45.6 11.8 16.1 11.4 38.8 20.9 11.5
RLET 38.3 9.1 11.5 7.1 34.2 12.1 6.9
NLProofS 43.2 8.2 11.2 6.9 42.9 17.3 6.9
FAME 43.4 13.8 16.6 12.4 40.6 19.9 11.9
GPT4-CoT 44.1 12.1 15.4 10.8 43.1 20.6 10.8
GPT4-ToT 43.3 12.0 15.8 11.0 43.9 20.0 11.0
GPT4-ReAct 45.8 12.9 14.1 10.5 43.5 21.5 10.5
Seer (Ours) 47.1 13.8 17.4 12.9 45.1 18.8 12.9
Table 2: Experiment results on EntailmentBank. Bold and underlined texts highlight the best method and the runner-up. RLET is based on DeBERTa-large (He et al., 2023), while all other methods are based on T5-large. All baseline results come from published papers. We use the gpt-4-1106-preview version for GPT-4.
5 Result Analysis
5.1 Structured Reasoning
EntailmentBank
As shown in Table 2, our Seer outperforms all baseline methods on the most strict metric, "Overall AllCorrect", across all three tasks. Specifically, our method achieved an absolute improvement of 1.7%/1.4%/1.0% in Task 1/2/3 compared to the strongest baseline. The steps dimension, i.e., premises selection, is the core of EntailmentBank [^3] A comprehensive error analysis is detailed in Appendix G., contributing to enhancing the accuracy of both leaves and intermediates dimensions, thereby improving the Overall AllCorrect metric. (1) Compared to SOTA supervised methods, such as NLProofs and FAME, our method exhibits significant advantages in the steps dimension. This demonstrates that focusing solely on isolated single-step reasoning through supervised learning may yield suboptimal solutions in intricate structured reasoning tasks, even though employing advanced planning algorithms, such as Monte-Carlo planning in FAME. (2) Compared to SOTA RL-based method, our method outperforms RLET by absolute 5.8%/9.0%/6.0% in Task 1/2/3. Our method employs a generative model as the policy to circumvent the issue of enumerating actions, facilitating the policy’s understanding of structured reasoning tasks (generating potential actions by itself). Moreover, our proposed structure-based return more effectively captures the tree-structured logical dependencies between steps and is able to assign stable returns for equivalent trajectories, which significantly improves reasoning abilities. Subsequent ablation studies will further demonstrate this. (3) Compared to GPT-4 with CoT, ToT, and ReAct, our method achieves an absolute improvement of 1.9% in Task 3. Although GPT-4 exhibits outstanding reasoning capabilities surpassing many baselines, its performance relies on a vast number of parameters. Detailed prompt examples for GPT-4 are provided in the Appendix F.
Method Task 1 Task 2
SI+Halter 72.4 55.9
SI+Halter+Search 83.2 72.9
FAME 91.5 78.2
Seer (Ours) 92.7 85.6
Table 3: Experiment results on the EntailmentBankQA. SI is based on Chinchilla-7B (Hoffmann et al., 2022).
EntailmentBankQA
Following (Creswell and Shanahan, 2022), we introduce the halter module to generate answers based on $T_{\text{pred}}$ and substitute hypothesis with question and option during the reasoning process. As illustrated in Table 3, our method surpasses FAME by an absolute margin of 1.2%/7.4% in Task 1/2. While both FAME and SI are supervised methods, FAME significantly outperforms SI by enhancing the model’s reasoning and exploration capabilities through Monte-Carlo planning. However, our method enhances the structured reasoning capabilities of the policy rather than focusing solely on single-step reasoning, which can significantly improve the quality of the entailment tree to aid in answering, especially in complex reasoning environments.
Method SCONE GSM8K AQUA-RAT AR-LSAT
Answer Accuracy
STREET 69.6 10.4 28.7 28.0
GPT4 † 66.0 94.0 78.0 32.0
SEER (Ours) 72.4 21.4 37.6 33.5
Reasoning Graph Accuracy
STREET 60.0 0.7 0.0 0.0
GPT4 † 32.0 10.0 4.0 2.0
SEER (Ours) 64.8 13.4 8.1 7.2
Table 4: Experiment results on STREET benchmark. † indicates we recorded the best results in CoT, ReAct, and ToT for brevity.
STREET
As shown in Table 4, our method significantly outperforms the SOTA baselines in graph-structured reasoning across all datasets. While GPT-4 is adept at answering questions, it suffers from hallucinations (Rawte et al., 2023), which leads to performance degradation in structured reasoning and explanations. Since SCONE contains sufficient data as well as similar QA and reasoning patterns, we observe that the STREET method would outperform GPT-4 on SCONE. However, our Seer further enhances performance by bolstering structured reasoning capabilities. The Reasoning Graph Accuracy is a very strict metric (Ribeiro et al., 2023). Nevertheless, our method achieved absolute improvements of 4.8%/3.4%/4.1%/5.2% on various datasets, respectively. By obtaining high-quality reasoning graphs, our method significantly improves trustworthiness. In reasoning graphs, a state may have multiple parent nodes. However, our structure-based return (equation 6) still precisely describes the expected cumulative rewards for each state, thereby facilitating reasoning performance in graph-structure reasoning.
5.2 Cross-dataset Performance
Method eQASC eOBQA
P@1 NDCG P@1 NDCG
EntailmentWriter 52.48 73.14 69.07 89.05
EntailmentWriter-Iter 52.56 73.28 72.15 90.19
METGEN 55.81 74.19 74.89 90.50
FAME 53.36 79.64 73.09 89.32
GPT-4 54.00 88.82 85.36 91.19
Seer (Ours) 60.33 89.76 77.50 94.62
Table 5: Cross-dataset performance on the eQASC and eOBQA.
To evaluate the generalization performance, we conduct cross-dataset experiments on eQASC and eOBQA (Jhamtani and Clark, 2020). More details can be found in Appendix B. We apply the policy of Task 2 for selection without training on eQASC or eOBQA. As illustrated in Table 5, our method exhibits significant superiority in cross-dataset generalization. Compared to supervised methods, our Seer, following the inherent structural nature of entailment trees, can better capture the logical dependencies between reasoning steps, which can effectively promote the generalization ability of the policy. The experimental results further validate the effectiveness of our method.
5.3 Ablation Studies
Method Leaves Steps Intermediates Overall
Seer (Ours) 13.8 12.9 18.8 12.9
w/o redundant 13.2 12.6 18.5 12.3
w/o structure-based return 12.9 11.7 18.5 11.1
w/o RL 11.1 10.2 17.3 10.0
Table 6: Ablation study of each component.
Refer to caption (a) $r_{\text{redundant}}$ Refer to caption (b) $\beta$
Figure 3: Parameter sensitivity analysis.
To evaluate the contribution of each component, we conduct extensive ablation studies. As shown in Table 6, we investigate three different variations of Seer in Task 3: (1) w/o redundant neglects redundant steps by assigning a reward of -1. (2) w/o structure-based return removes the structure-based return and calculates it in a chained trajectory (equation 5). (3) w/o RL removes the RL phase, relying solely on supervised warm-up. We discovered that overlooking redundant steps may potentially inhibit the exploration of policy, leading to a performance decline. Then, the results in Table 6 demonstrate that removing the structure-based return severely affects the performance. It not only adequately addresses the equivalent trajectory problems, but also elegantly captures the logical relationships inherent in entailment trees, which is crucial for tree-structured reasoning. Furthermore, the results show that removing the RL phase reduces performance by 2.9% of Overall Allcorrect, which is a significant impact for this strict metric. This demonstrates that relying solely on supervised learning may overlook the logical relationships in structured reasoning, thereby falling into suboptimal solutions.
5.4 Parameter Sensitivity Analysis
As illustrated in Figure 3, we further investigate the impact of $r_{\text{redundant}}$ and $\beta$ on the performance in Task 3. We observe that compared to treating redundant and erroneous steps equally ($r_{\text{redundant}}=-1$), not penalizing ($r_{\text{redundant}}=0$) may have more detrimental effects, which allows for unrestricted exploration. Moreover, a suitable $\beta$ (the coefficient of entropy bonus) is crucial for performance enhancement, as it encourages the policy to break away from the "stereotypes" of supervised warm-up.
6 Conclusions
We propose Seer, a novel approach that facilitates structured reasoning and explanation via RL. To our knowledge, Seer is the first general framework capable of enhancing chained, tree-based, and graph-based structured reasoning. Our structure-based return precisely delineates the hierarchical and branching structure inherent in structured reasoning, effectively facilitating reasoning ability. Furthermore, Seer employs a generative model as the policy and refines the reward function, ingeniously circumventing the limitations of existing works. Comprehensive experimental results demonstrate that Seer significantly outperforms state-of-the-art methods and exhibits outstanding cross-dataset generalization performance.
Appendix A Algorithm Details
Input: Structured reasoning dataset $\mathcal{D}$; Training epochs $N_{\text{warmup}}$, $N$ and $N_{K}$; batch size $b_{\text{warmup}}$ and $b_{\text{mini}}$;
Output: The optimal parameter of policy
/* (1) Supervised Warm-up phase */
1 Initialise policy parameters $\pi_{\theta}$
2 Convert $\mathcal{D}$ into single-step data $\mathcal{D}_{step}$
3 for epoch $=1$ to $N_{\text{warmup}}$ do
4       for i = 1 … $|\mathcal{D}_{step}|/b_{\text{warmup}}$ do
5             sample minibatch from $\mathcal{D}_{\text{step}}$
6             update parameters $\pi_{\theta}$ by Eq. 11
7
/* (2) RL phase */
8 Initialise the critic parameters $V$
9 for epoch $=1$ to $N$ do
10       Initialise training buffer $\mathcal{B}\leftarrow\varnothing$
       // Filling the replay buffer
11       while $\mathcal{B}$ not full do
12             sample $\{h,X,T_{\text{gold}}\}$ from $\mathcal{D}$
13             collect trajectory $\tau$ via $\pi_{\theta}$
14             fill buffer $\mathcal{B}\leftarrow\{s_{t},a_{t},r_{t}\}$ from $\tau$
15
      // Performing k-epoch updates per buffer
[…]
17             sample $\{(s_{t},a_{t},r_{t})\}_{b_{\text{mini}}}$ from $\mathcal{B}$
18             compute $\mathcal{E}(\pi^{\prime}_{\theta})$ and $\hat{A}_{t}$ by Eq. 8,  9
19             update policy $\pi_{\theta}$ by Eq. 7
20             update critic $V$ by Eq. 10
21
22
Algorithm 1 The training process of Seer
Algorithm 1 provides a detailed description of the overall training process for Seer, which primarily consists of two phases: supervised warm-up and reinforcement learning. In the supervised warm-up phase, it initially decomposes structured reasoning into single-step reasoning data (Line 2). Then, it employs supervised learning to guide the policy $\pi_{\theta}$ to quickly adapt to the complex reasoning environments (Lines 3-6). This is particularly beneficial when the number of parameters in the policy is relatively small (Akyurek et al., 2023; Liu et al., 2023). In the RL phase, it initially populates the replay buffer $\mathcal{B}$ through the policy $\pi_{\theta}$ (Lines 9-13). Then, it performs $N_{K}$ updates for each replay buffer to improve sample efficiency (Lines 14-18).
Appendix B Datasets Details
Task Name Task Domain # Questions # Reasoning Steps Reasoning Type Answer Type
EntailmentBank Science 1,840 5,881 Tree-structured /
EntailmentBankQA (ARC) Science 1,840 5,881 Tree-structured 4-Way MC
SCONE Processes 14,574 130,482 Graph-structured State Pred.
GSM8K Math 1,030 4,666 Graph-structured Number
AQUA-RAT Math 1,152 7,179 Graph-structured 5-Way MC
AR-LSAT Logic 500 2,885 Graph-structured 5-Way MC
Table 7: Datasets Statistics of Structured Reasoning.
Datasets of Structured Reasoning
Table 7 describes the statistics of datasets in detail. In the answer types, “K-Way MC” stands for multiple choice answer with K options.
EntailmentBank (Dalvi et al., 2021) comprises 1,840 expert-annotated entailment trees with an average of 7.6 nodes spanning across 3.2 entailment steps. The facts are derived from the WorldTree V2 corpus (Xie et al., 2020). Following Dalvi et al. (2021), based on different facts $X$, there are three progressively more challenging tasks: Task1(no-distractor), Task2(distractor), and Task3(full-corpus). For GPT-4, we employ all the data in Task3 from EntailmentBank to evaluate its performance. As EntailmentBank was initially designed for post-hoc tree reconstruction instead of QA, Tafjord et al. (2022) convert it to choose the correct answer given the multiple choice options instead of deriving hypothesis $h$, namely EntailmentBankQA.
In the STREET benchmark, Ribeiro et al. (2023) converts each QA dataset, such as ARC (Clark et al., 2018), SCONE (Long et al., 2016), GSM8K (Cobbe et al., 2021), AQUA-RAT (Ling et al., 2017) and AR-LSAT (Zhong et al., 2021), to graph-structured explanation formats and needs to answer the question base on the predicted reasoning graphs. Please note that as ARC in STREET is congruent with Task 1 of EntailmentBankQA (Ribeiro et al., 2023), we do not repeat the experiment for this task in Table 4. Due to the high cost of GPT-4, we randomly sample 50 instances from each dataset in the STREET benchmark as an evaluation of GPT-4’s performance.
Datasets of Cross-dataset Experiments
To evaluate the generalization performance of our method, following (Hong et al., 2022), we conduct cross-dataset experiments on eQASC and eOBQA (Jhamtani and Clark, 2020), which collect one-step entailment trees for questions from QASC (Khot et al., 2020) and OpenBookQA (Mihaylov et al., 2018), respectively. The goal of this task is to select valid one-step trees from the candidate set and evaluate the results with P@1 and NDCG metrics (Hong et al., 2022). Questions with no valid tree are filtered. The candidate sets for eQASC and eOBQA are composed of 10 and 3 sentences respectively.
Appendix C Implementation Details
C.1 Stopping condition
For a fair comparison, we use T5-large model as the policy. However, we observe that due to the number of parameters as well as the issue about only a few "End" instances, T5-large tends to perform "Reason" actions more frequently. Moreover, unlike GPT-4, T5-large is less able to recognize when a hypothesis has been inferred and when to stop. Therefore, we attach other stopping conditions in addition to the "End" action: (1) The semantic similarity between the intermediate conclusion and the hypothesis exceeds a predefined threshold, i.e., BLEURT($i_{*}$, $h$) > 1. (2) Exceeding the maximum number of steps (set to 20 in this paper).
C.2 Alignment algorithm
Following (Dalvi et al., 2021), we evaluate the intermediate steps based on Jaccard similarity. Specifically, the intermediate nodes $i_{*}$ in $T_{\text{pred}}$ are aligned with the intermediate nodes in $T_{\text{gold}}$ that have the maximum Jaccard similarity. If the Jaccard similarity between the intermediate node in $T_{\text{pred}}$ and all intermediate nodes in $T_{\text{gold}}$ is zero, it is aligned with "NULL". Note that only the intermediate node that is perfectly matched, i.e., the Jaccard similarity is 1, is considered as a correct step. Figure 5 provides a detailed illustration of this process. The alignment process is similar in the reasoning graphs (Ribeiro et al., 2023).
C.3 Retriever for Task 3
In Task 3 of EntailmentBank, it is first necessary to retrieve relevant sentences from the corpus (Dalvi et al., 2021). The focus of this paper is to enhance the structured reasoning ability of the policy. Therefore, we adopt the retriever and its associated parameters from previous work (Hong et al., 2023), which is based on the all-mpnet-base-v2 model (Reimers and Gurevych, 2019). For a fair comparison, we retrieve the top 25 most relevant sentences as $X$ for Task 3.
C.4 Entailment Module
The entailment module is also based on T5-large, taking premises as input and generating intermediate conclusions. Our primary focus is to enhance the structured reasoning ability of the policy through a reasonable RL method, such as, our proposed Seer in this paper. Therefore, similar to the retriever, we employ the entailment module that has already been trained in previous work (Hong et al., 2023), which also aids in a fair comparison.
C.5 Halter Module
In EntailmentBankQA, we employ the Halter module (Creswell and Shanahan, 2022) to answer questions based on the predicted entailment trees. In this paper, the Halter module is built upon the T5-large model. The module is trained with a learning rate of 1e-5 and a batch size of 16.
C.6 Entailment Tree Construction
To evaluate the correctness of each step, we have to reconstruct the trajectory into an entailment tree $T_{\text{pred}}$ for comparison with $T_{\text{gold}}$. Figure 5 illustrates this process. We consider the last intermediate conclusion as the hypothesis and then construct the predicted entailment tree based on the reasoning relationship of each step (each step is equivalent to a subtree). The reconstruction process is similar in the reasoning graphs (Ribeiro et al., 2023).
C.7 Running time
In our experimental environment, the average training time per entailment tree for our Seer is 6.98 seconds, and the average inference time per entailment tree for Seer is 3.91 seconds. The inference time per entailment tree for RLET (Liu et al., 2022) is 9.34 seconds. The inference time per entailment tree for FAME (Hong et al., 2023) is 30.77 seconds, as reported in their papers. FAME leverages Monte-Carlo planning, necessitating the exploration of numerous nodes to enhance the reasoning capability of the policy, thus requiring considerable computational time. Our proposed Seer significantly surpasses FAME in terms of both efficiency and effectiveness, representing substantial progress.
C.8 Experiment Environments
All experiments were conducted on Ubuntu 22.04 equipped with NVIDIA A100 GPUs. Our code mainly depends on python 3.10444https://www.python.org/ and PyTorch 2.0.1555https://pytorch.org/. The pre-trained language models are derived from HuggingFace Transformers [^6] https://huggingface.co/.
Appendix D Metrics Details
For EntailmentBank, we follow (Dalvi et al., 2021) and evaluate the entailment tree $T_{\text{pred}}$ with three dimensions:
$\bullet$ Leaves: To evaluate the leaf nodes of $T_{\text{pred}}$, we compute F1 by comparing $X_{\text{pred}}$ to $X_{\text{gold}}$.
$\bullet$ Steps: To evaluate the structural correctness of each step, we compare all steps between $T_{\text{pred}}$ and $T_{\text{gold}}$ and compute F1. A predicted step is considered structurally correct if its premises (e.g., $x_{*}$, $i_{*}$) exactly match the gold premises.
$\bullet$ Intermediates: To evaluate the intermediate conclusions, we compare the aligned intermediate conclusions and compute F1. A predicted intermediate conclusion is deemed correct if the BLEURT score (Sellam et al., 2020) exceeds 0.28.
For each dimension, AllCorrect score is 1 if F1 is 1, otherwise 0. Given the above scores, we employ the Overall AllCorrect metric to comprehensively evaluate $T_{\text{pred}}$, which takes a value of 1 if and only if all leaves, steps, and intermediates are all correct. Note that this is an extremely strict metric, where any deviations in $T_{\text{pred}}$ will result in a score of 0.
For the STREET benchmark, we follow (Ribeiro et al., 2023) to evaluate in terms of both the answer to the question and the quality of the reasoning graphs.
$\bullet$ Answer Accuracy: This metric measures the ability to correctly answer questions. The answer accuracy serves as an upper bound for other metrics, as any reasoning graph generated with incorrect answers is also labeled as incorrect.
$\bullet$ Reasoning Graph Accuracy: This metric compares the predicted reasoning graph and the golden reasoning graph from the aspects of the graph structure and the content of intermediate conclusions. Please note that this is a stringent metric, with minor deviations from the golden reasoning graph resulting in the prediction incorrect.
Appendix E Illustrations and Case Study of Seer
Given a hypothesis $h$ and initial facts $X$, we first obtain the trajectory through the resoning process, as shown in Figure 4. The state update follows the markov decision process, meaning the current state only depends on the previous state. Figure 4 is an erroneous reasoning example to better illustrate the following steps. Then, we convert the trajectory $\tau$ into an entailment tree $T_{\text{pred}}$ and align it with $T_{\text{gold}}$ to assign reward for each intermediate conclusion (Figure 5). Furthermore, Figure 6 elucidates the issue of equivalent trajectories, and previous work can not accurately describe the logical relationship between different states in entailment trees.
Figure 4: An illustration of the reasoning process of Seer. We start from the initial state $s_{1}$ where existing entailment steps $P_{1}=\varnothing$ and candidate sentences $C_{1}=X$. In each step, we sample an action and update the state until the reasoning is done. For the "Reason" action, we sent the premises to the entailment module. The new conclusion is added to the $C$, the premises is removed from $C$ and the entailment step is added to the $P$. For the "End" action, we end the reasoning process and output the trajectory.
Figure 5: An illustration of the reward and alignment process of Seer. Each reasoning step is actually a subtree (the same in the reasoning graph (Ribeiro et al., 2023)). (1) We construct $T_{\text{pred}}$ using the last intermediate conclusion ($i_{4}$ in this example) as the hypothesis. (2) We calculate the Jaccard similarity between the intermediate node ($i_{*}$) in $T_{\text{pred}}$ and each golden intermediate node in $T_{\text{gold}}$ ($\hat{i}_{1}$ and $h$ in this example), and align with the maximum Jaccard similarity. In this example, $i_{1}$ is aligned with $\hat{i}_{1}$ due to $\text{JS}(i_{1},\hat{i}_{1})=1$. $i_{2}$ is aligned with "NULL". $i_{4}$ is aligned with $\hat{i}_{1}$ due to $\text{JS}(i_{4},\hat{i}_{1})=0.5$ and $\text{JS}(i_{4},h)=0.4$. (3) We assign rewards based on the alignment results. Note that $i_{3}$ ($s_{3}$) is a redundant step. $r_{1}=1$, $r_{2}=-1$, $r_{3}=-0.5$, and $r_{4}=-1$. Evidently, the reward for each state originates from the tree structure rather than chained trajectory. Therefore, the return of each state should also follow the tree structure (or graph structure in reasoning graphs) rather than chained trajectory.
Figure 6: An illustration of the equivalent trajectory and the definition of return. As the reasoning steps of $s_{1}$, $s_{2}$, and $s_{3}$ are mutually independent, the execution order among these steps can be arbitrary. Thus, $\tau$ and $\tau^{\prime}$ are equivalent trajectories because they can be converted to the same entailment tree (Dalvi et al., 2021). As shown in blue box, previous work (Liu et al., 2022) defines the return (a.k.a expected cumulative reward) in a chained trajectory and would assign different returns to $s_{1}$ and $s_{2}$ in these equivalent trajectories. In contrast, as shown in red box, our structure-based return is defined based on the tree or graph structure inherent in structured reasoning, which is the same source of rewards. Our structure-based return will consistently allocate stable returns to equivalent trajectories, thereby promoting training stability and convergence. Furthermore, maintaining consistency between the sources of rewards and returns can significantly enhance the effectiveness of the policy.
Appendix F Prompts for GPT-4
Figure 7: A Chain-of-Thought prompt for GPT-4.
Figure 8: A ReAct prompt for GPT-4. "Thought" and "Action" query GPT-4 in two rounds.
Figure 9: A Tree of Thought prompt (Thought Generator) for GPT-4.
Figure 10: A Tree of Thought prompt (State Evaluator) for GPT-4.
Figures 7 and 8 show the CoT (Wei et al., 2022) and ReAct (Yao et al., 2023b) prompts for GPT-4, and figures 9 and 10 show the prompts of thought generator and state evaluator in ToT (Yao et al., 2023a), respectively. We provide a detailed introduction to the task definition and guide the model to respond in the required format. We randomly selected three examples for in context learning. For a fair comparison, we use the same examples for CoT and ReAct, attributing similar thought to them. ReAct divides the dialogue into two rounds, "Thought" and "Action", to query GPT-4. For ToT, following (Yao et al., 2023a), we generate candidate actions using a thought generator and subsequently select and execute the optimal action through a state evaluator. Due to its exceptional reasoning capabilities and self-evaluation strategy, ToT achieves superior results compared to CoT and ReAct, as demonstrated in Table 2. However, ToT requires higher costs in comparison to CoT and ReAct.
Appendix G Error Analysis
We conduct comprehensive error analysis on Task2 and Task3 of EntailmentBank.
G.1 Error Analysis of Task2
We randomly sample 50 entailment trees where Seer made incorrect reasoning. We find the following four types of errors.
(1) Reasoning Step Error (62%). This is the main source of errors, and predominantly depends on whether the policy can select the correct premise. We observe that a small portion of the errors (accounting for 12.9% of this error type) use all the gold leaves, but have errors in the combination order. Compared to other reasoning step errors, the policy identified the correct premise. For example, the gold steps is "$x_{24}\;\&\;x_{5}\rightarrow i_{1}$; $i_{1}\;\&\;x_{23}\rightarrow h$" and the error steps predicted by Seer is "$x_{23}\;\&\;x_{5}\rightarrow i_{1}$; $i_{1}\;\&\;x_{24}\rightarrow i_{2}$".
(2) Early Termination Error (18%). We observe that the reasoning process may terminate prematurely and the existing entailment steps are all correct. On one hand, T5-large outputs ’End’ prematurely, unlike GPT-4 which can accurately judge when to stop. On the other hand, the entailment module might erroneously infer a hypothesis, leading to premature termination.
(3) Intermediate Conclusion Error (10%). This error type is different from the above error (where the entailment module prematurely infers a hypothesis). Intermediate conclusion error denotes errors triggered by incorrect entailment in the intermediate conclusions, despite having both correct leaves and steps. For a fair comparison, we used the entailment module that has already been trained in previous work Hong et al. (2023). These errors can be mitigated by training a better entailment module. Furthermore, these errors are correct in the reasoning part, which is also the our focus in this paper.
(4) Imperfect Evaluation (10%). We discover that some trees deemed as invalid are in fact valid, indicating that current automated metrics underestimate the validity of the trees. The most common reason is that there are multiple valid ways to construct an entailment tree. For example, consider the structure of a gold tree: "$x_{1}\;\&\;x_{2}\;\&\;x_{3}\rightarrow h$" may be predicted as: "$x_{1}\;\&\;x_{2}\rightarrow i_{1}$; $i_{1}\;\&\;x_{3}\rightarrow i_{2}$".
G.2 Error Analysis of Task3
Task 3 requires retrieving an initial set of facts $X$ from the corpus. Therefore, in addition to the errors in Task 2 described above, we found that Task 3 has its own unique set of errors.
(1) Missing Gold Leaves Error (58%). Missing gold leaves error refers to the situation where the gold leaves are not included in the facts $X$ retrieved from the corpus. This situation will inevitably lead to an error in the predicted entailment tree, regardless of how powerful the policy is. The bottleneck of this error lies in the retrieval model. For a fair comparison, we directly use the retrieval model that has already been trained in previous work (Hong et al., 2023).
(2) Reasoning Step Error (42%). This error is similar to those in Task 2.
Here we will not discuss other errors included in Task 2, as we found that a major bottleneck of Task 3 is how to retrieve more gold leaves. The focus of this paper is to enhance structured reasoning through a sensible RL approach. In future work, we will continue to consider how to integrate the retrieval module into RL, further improving performance in Task 3.